How many grams of solid potassium hypochlorite should be added to 1.50 L of an 8.99×10-2 M hypochlorous acid solution to prepare a buffer with a pH of 8.453?

Grams potassium hypochlorite=
Answer #1

pKa of hypochlorous acid = 7.53

pH = pKa + log ( salt / acid )

Moles of hypochlorous acid = Molarity * volume

= 8.99*10^-2 * 1.50

= 0.13485

8.453 = 7.53 + log ( salt / 0.13485 )

Salt = 1.13

Mass = mole * molar mass

= 1.13 * 90.55

= 102.3215 grams

If you have any query please comment

If you satisfied with the solution please like it thankxx